# Optimal quantum simulation of linear non-unitary dynamics

### v1

@@ 1 Introduction
Non-unitary dynamics appears in other exciting application areas, including dissipative non-linear dynamics by Carleman linearization [LøKK+21, LAF+23], partial differential equations [CLO21], and linear matrix equations [SLBB25], but exploration is hindered by a more limited understanding of how to best approach Equation 1. Within the context of quantum algorithms, the goal is to prepare a quantum state that approximates the normalized $\lvert\vec{u}(t)\rangle\doteq\vec{u}(t)/|\vec{u}(t)|$ for some $t>0$, which has the formal solution
@@
in terms of the time-ordering operator $\mathcal{T}$. In the special case of anti-Hermitian $A(t)$ and a zero forcing term $\vec{b}(t)=0$, Equation 1 is the famous Schrödinger equation that underlies quantum mechanics and the Hamiltonian simulation problem. In the quantum query setting, Equation 2 is solved by queries to a unitary $\textsc{Prep}_{\lvert\vec{u}_{0}\rangle}$ that prepares the initial state $\lvert\vec{u}_{0}\rangle$ from a computational basis state, and a unitary Definition 1 that block-encodes the operator $A$ (see Definition 14 for the time-dependent case), which is a well-established access model that generalizes sparse matrices with efficiently computable entries [ATS03].
@@
[BC24] Dyson series+QLSP $A+A^{\dagger}\succeq 0$ $u_{\text{r}}\alpha_{A}t\log\frac{\lambda t}{\epsilon}\log\frac{1}{\epsilon}$ $u_{\text{r}}\alpha_{A}t\log\frac{1}{\epsilon}$
@@
In Table 1, we compare quantum algorithms for the general case of non-Hermitian $A$ with the results of this work. To compare algorithmic advances, we focus on the homogeneous time-independent case where $\vec{u}(t)=e^{-At}\lvert\vec{u}_{0}\rangle$. These algorithms require some ‘stability’ condition on $A$, such as non-negative eigenvalues $\mathrm{Re}[\Lambda]\succeq 0$ [Ber14, BCOW17, CL20, LS24a] in the eigendecomposition $A=S\Lambda S^{-1}$. In this setting, all known results are suboptimal by some logarithmic factor of time $t$, error $\epsilon$, or condition number $\kappa=\|S\|\|S^{-1}\|$, which is a measure of matrix non-normality. Under the stronger condition where A-A has non-positive logarithmic norm or $L\doteq\frac{1}{2}(A+A^{\dagger})\succeq 0$ [BC24], the recent Linear-Combination-of-Hamiltonian (LCHS) [ALL23, ACL23] approach achieves better scaling, in particular optimal scaling in $u_{\text{lchs}}\doteq|\vec{u}_{0}|/|\vec{u}(t)|$ queries to $\textsc{Prep}_{\lvert\vec{u}_{0}\rangle}$. However, the query complexity to $\textsc{Be}[A/\alpha_{A}]$ scales as $\mathcal{O}(\alpha_{A}t\log^{1+\mathcal{O}(1)}\frac{1}{\epsilon})$, which is suboptimal in error. Furthermore, a no-go result [ACL23, Proposition 7] suggested that achieving the optimal scaling $\mathcal{O}(\alpha_{A}t\log\frac{1}{\epsilon})$ was impossible within the existing LCHS framework. Moreover, known lower bounds are additive in each parameter, in contrast to the multiplicative upper bound. We note that scaling with other state-dependent parameters is possible, but these alternate approaches [Kro23, FLT23, LS24a] are incomparable and out-of-scope for this work.
@@
under some additional conditions on $\hat{f}$. Hence, one block-encodes the non-unitary time-evolution operator $\textsc{Be}[e^{-At}/\alpha_{\hat{f}}$] to additive error $\epsilon$ with normalization $\alpha_{\hat{f}}\doteq\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}|\hat{f}(k)|\mathrm{d}k$ by the Linear-Combination-of-Unitaries (LCU) lemma [CW12] after discretizing the integral in Equation 4 and truncating the discrete sum to finite $R$. The query complexity is determined by that of Hamiltonian simulation of the ‘most expensive’ unitary $e^{-i(\pm RL+H)t}$. As optimal Hamiltonian simulation in the large time limit requires $\alpha t+\mathcal{O}(\log\frac{1}{\epsilon})$ queries to the block-encoded Hamiltonian [LC17b, LC19, BMPW24], this has query complexity $\mathcal{O}(\alpha_{A}Rt+\log\frac{1}{\epsilon})$. If preparing $\lvert u(t)\rangle$ is desired, this is further multiplied by at least $\alpha_{\hat{f}}$ to account for amplitude amplification. Previously, a LCHS based on $\hat{f}_{\text{ACL23}}(k)\propto\big{(}(1-ik)e^{(1+ik)^{\varphi}}\big{)}^{-1}$ for any constant $\varphi\in(0,1)$ was shown to achieve a truncation threshold $R=\mathcal{O}(\log^{1/\varphi}\frac{1}{\epsilon})$ [ACL23]. Unfortunately, the optimal scaling of $\alpha_{\hat{f}}R=\mathcal{O}(\log\frac{1}{\epsilon})$ was unattainable as $\lim_{\varphi\rightarrow 1}\alpha_{\hat{f}_{\text{ACL23}}}=\infty$, and a no-go theorem [ACL23] forbids this scaling for any LCHS that satisfies Equation 4.
In this work, we resolve these limitations and present a query-optimal algorithm for this problem. Our key contributions are: 1) A generalized LCHS framework that achieves the optimal query complexity $\alpha_{\hat{f}}R=\mathcal{O}(\log\frac{1}{\epsilon})$ of block-encoding $e^{-At}$, circumventing a no-go result that applied to previous LCHS methods, and as shown in Figure 1, with a significant benefit in practical applications; 2) A novel, analytically simple, family of kernel functions that only approximate exponential decay. This is enabled by the generalized LCHS framework and is a conceptual departure from prior work that required an exact match; 3) A greatly simplified and more efficient implementation of the LCHS by proving that a simple, uniform discretization of the LCHS integral converges exponentially fast, leading to simpler quantum circuits than previous non-uniform quadrature methods; 4) New provable and empirical lower bounds on the query complexity, showing that our algorithm’s performance is optimal up to a small constant factor (which we estimate to be approximately three), making further significant improvements in this framework infeasible. These insights are formalized in our main results, which we now summarize.
@@
This is formalized in our main mathematical result Theorem 1, stated for general time-dependent $A$, which establishes rigorous error bounds on Equation 5. A consequence of the exact exponential decay equality Equation 4 is that previous approaches required $\hat{f}(z)$ to decay uniformly to zero in the lower complex plane [ALL23, ACL23], which is also a key assumption of the previous no-go. In contrast, we only require decay $\lim_{|z|\rightarrow\infty}\hat{f}(z)=0$ in an infinite strip of finite width, which implies that $f(x)$ only approximates exponential decay. This flexibility allows us to study a new $4$-parameter family of kernel functions
@@
We prove the optimal cost scaling in our main algorithmic result Theorem 2 by instantiating Theorem 1 with the explicit choice $\hat{f}_{2}(x;\gamma,c)\doteq\hat{f}_{2,1}(x;\gamma,c)$ illustrated in Figure 2 and alternatively defined in terms of elementary functions as follows:
@@
where $\mathrm{erfc}$ is the complementary error function, $c$ is any positive constant, $\gamma=\mathcal{O}(\frac{1}{c}\log^{1/2}\frac{1}{\epsilon})$, and $R=\mathcal{O}(c\gamma^{2})$. We design $f_{2}$ to be analytic in the entire complex plane, which is motivated by the super-polynomial to super-exponential decay of entire functions in the Fourier domain. $f_{2}$ is also the convolution of $e^{-|x+c|+c}$ with a Gaussian kernel, and its simple Fourier transform and constant normalization satisfy the conditions of Theorem 1 and allows us to derive simple and reasonably tight upper bounds on the performance of our LCHS. Although infinitely many kernel functions satisfy Theorem 1, not all lead to an optimal-scaling LCHS, even among those where $f$ is entire like $f_{2}$. For instance, the approximations $f_{\text{HMS22}}=\frac{1}{2}e^{-x}\mathrm{erfc}(\gamma-x)$ [HMS+22] and $f_{1}(x;\gamma)$ both turn out to have unbounded norm $\alpha_{\hat{f}}\rightarrow\infty$ when $\gamma\rightarrow\infty$.
Figure 2: Our entire approximation $f_{2}$ to exponential decay converges to $\lim_{\gamma\rightarrow\infty}\hat{f}_{2}(k,\gamma,c)=e^{-|x+c|+c}$.
@@
For any $t\geq 0$, let $A=L+iH$ be as described in Theorem 1. For any $\epsilon_{\text{lchs}}\in(0,0.9027$...] and any $c>0$, the LCHS of time-evolution generated by $kL+H$ and the kernel function $\hat{f}_{2}\doteq\hat{f}_{2,1}$ from  Equation 6 with $\gamma=\frac{1}{c}\sqrt{c+\log\frac{1+1/({2\pi})}{\epsilon_{\text{lchs}}}}$ and $R=2c\gamma^{2}$ satisfies
@@
We also greatly improve the practical implementation of block-encoding the LCHS. We prove in Theorem 3 that the error $\epsilon_{\text{quad}}$ of a uniform quadrature is exponentially convergent in the stepsize $h$ [TW14]. In contrast, previous analyses either showed polynomial error scaling for the uniform quadrature [ALL23], or used a complicated multi-step non-uniform Gauss-Legendre quadrature [ACL23] with still an asymptotically worse stepsize, possibly a larger block-encoding normalization with no simple closed-form expression due to non-uniform weights, a more complicated quantum circuit implementation for non-uniform quadrature points, and a complicated analysis involving higher derivatives.
@@
Choose the LCHS parameters $\epsilon_{\text{lchs}},R,\gamma,c$ according to Theorem 2. For any $\epsilon_{\text{quad}}\in(0,4/15]$, and any stepsize ${h}\leq{\pi}/\left(\frac{\|L\|_{L^{1}}}{2}+\log\left(\frac{64e^{3c/2}}{15\epsilon_{\text{quad}}}\right)\right)$ such that $R/h$ is an integer, let $I_{h}$ be the uniform discretization of $O_{R,\gamma,c}(t)$. Then
@@
We then block-encode non-unitary time-evolution in Theorem 4 to error $\epsilon$ by an explicit and efficient quantum circuit implementation of the Linear-Combination-of-Unitaries (LCU) method [CW12] applied to Theorems 2 and 3, for instance, with $\epsilon_{\text{quad}}=\epsilon_{\text{lchs}}=\epsilon/3$. Note that for any reasonably small error, the difference in block-encoding normalization $\alpha_{\hat{f}_{2},R,h}$ from $\alpha_{\hat{f}_{2}}$ is extremely small and contributes a multiplicative constant cost that is $1+\mathcal{O}(\epsilon)$. Our use of the uniform quadrature greatly simplifies the explicit quantum circuit implementation, which we provide. As is known in prior art, the query complexity of $\textsc{Sel}_{R}$ for selecting time-evolution for different $k_{j}$ in Theorem 4, even with the additional control by $\lvert j\rangle$, can be reduced to that of $U_{0}(\pm Rt)$ without any controls by any choice of Hamiltonian simulation algorithm. In the time-independent case, Hamiltonian simulation by quantum signal processing and qubitization [LC17b, LC19] leads to an overall $\mathcal{O}(R\alpha t+\log\frac{1}{\epsilon})=\mathcal{O}(\alpha t\log\frac{1}{\epsilon})$ query complexity.
@@
For any $\epsilon\in(0,4/5]$, and under the same condition as Theorems 2 and 3, we may block-encode an operator $I_{h}^{\prime}$ such that $\left\|U_{0}(t)-I_{h}^{\prime}\right\|\leq\epsilon$ by
@@
where Prep and $\overline{\textsc{Prep}}$ costs $\mathcal{O}\left(\log(\|L\|_{L^{1}}+\log\frac{1}{\epsilon})\log^{5/2}\frac{1}{\epsilon}\right)$ two-qubit gates, $\textsc{Sel}_{R}$ has the same query complexity $Q$ and gate complexity as simulating $\mathcal{T}e^{-i\int_{0}^{t}\pm RL(s)+H(s)\mathrm{d}s}$ plus $\mathcal{O}\big{(}Q(\log(\|L\|_{L^{1}}+\log\frac{1}{\epsilon})\big{)}$ two-qubit gates.
We also optimize the constant factors in the query complexity of our optimal LCHS. The query complexity proportionality constant ${\alpha_{\hat{f}}}\cdot R$ is loosely bounded by Theorem 2, such as by setting $c=1$ in Figure 1. Due to our use of the uniform quadrature, the LCU normalization factor $\alpha_{\hat{f},R}\approx\alpha_{\hat{f}}$ up to a vanishingly small error, unlike prior art [PJKW25] where $\alpha_{\hat{f},R}$ can be a significant constant factor larger than $\alpha_{\hat{f}}$ due to non-uniform quadrature weights. We obtain tighter upper bounds on approximation error by evaluating the integrals in Theorem 1 numerically, and then numerically minimizing the product ${\alpha_{\hat{f}_{2},R}}\cdot R$ for each target $\epsilon$ with respect to $c,\gamma,R$. As seen in Figure 1, the LCHS with the kernel function $\hat{f}_{2}$ strictly outperforms prior art [ACL23, PJKW25] in all parameter regimes, and further improvements of $\approx 25\%$ are obtained by also optimizing over $j,y$ in the more general family Equation 6. Moreover, the proofs of implementation details in Theorems 3 and 4 for the special case $\hat{f}_{2}$ generalize readily to $\hat{f}_{j,y}$.
@@
As $\alpha_{\hat{f}_{2}}=\Theta(1)$ is constant for any constant $c$, the query complexity of our LCHS to state preparation $\textsc{Prep}_{\lvert\vec{u}\rangle_{0}}$, similar to prior LCHS approaches, is optimal due to an $\Omega(u_{\text{lchs}})$ lower bound [ALWZ25, Corollary 8] based on quantum state discrimination. We note that existing lower bounds on $Q_{A}$ are quite loose. When $A$ is anti-Hermitian, this reduces to Hamiltonian simulation with $u_{\text{lchs}}=1$ and $Q_{A}=\Omega(\alpha_{A}t+\frac{\log 1/\epsilon}{\log\log 1/\epsilon})$ [BCK15]. For the general case of $A$ with $L\succeq 0$, it is only known that $Q_{A}=\Omega(\min((\alpha_{A}t)^{\alpha},\log^{\alpha}\frac{1}{\epsilon}))$ for some unknown parameter $\alpha\leq 1$ [ALWZ25]. In Theorem 5, we present a significantly tighter quantum search lower bound $Q_{A}=\Omega(u_{\text{lchs}}\alpha_{A}t)$, which proves optimality of our LCHS up to a logarithmic factor in $u_{\text{lchs}}$ when $\epsilon$ is constant.
@@
Let $p_{n}$ be the polynomial of minimum degree $n=\min_{f}\widetilde{\mathrm{deg}}_{\epsilon^{\prime}}(f)$ that satisfies Equation 16 for some constant $\alpha$ – for example, see Figure 3. By Quantum Singular Value Transformations (QSVT) [GSLW19], $f(A/\alpha_{A})$ is block-encoded with a normalization factor at most $2(1+\epsilon^{\prime})$ using exactly $\widetilde{\mathrm{deg}}_{\epsilon^{\prime}}(f)$ queries to $\textsc{Be}[A/\alpha_{A}]$, which is optimal including constant factors by a matching $\widetilde{\mathrm{deg}}_{\epsilon^{\prime}}(f)$ [MS24, Lan25] lower bound. Hence, our LCHS can be shown to be asymptotically optimal in all parameters if $\min_{f}\widetilde{\mathrm{deg}}_{\epsilon^{\prime}}(f)=\Omega(\tau\log\frac{1}{\epsilon})$, where minimization is over all $f$ satisfying Equation 16. We compute $p_{n}$ by writing the optimization for $\epsilon$ over $p_{n}$ satisfying Equation 16 as a semi-infinite linear program [HK93] and in Figure 3 observe an empirical scaling of $\min_{f}\widetilde{\mathrm{deg}}_{\epsilon^{\prime}}(f)\approx 0.31\tau\log\frac{e}{31\epsilon}$ when $\alpha=e$. To our knowledge, this is the first computation of the optimal bounded polynomial for $e^{-\tau x}$ for $\tau x\geq 0$ and is of independent interest. Based on these empirical fits, our optimized LCHS $\hat{f}_{2}$ and $\hat{f}_{j,y}$ cannot be improved by more than a factor of $4$ and $3$ respectively.
@@
The rest of the manuscript is organized as follows. In Section 1.1, we discuss some possible extensions of our results and conclude with some open problems. In Section 2, we prove our main mathematical result Theorem 1 that a LCHS can be constructed from a kernel function that only approximates exponential decay. In Section 3, we prove our main algorithmic result Theorem 2 that a LCHS based on the kernel function $\hat{f}_{2}$ block-encodes $U_{0}(t)$ with the desired scaling in normalization and truncation radius of $\alpha_{\hat{f}_{2}}R=\mathcal{O}(\log\frac{1}{\epsilon})$. In Section 4, we prove Theorem 3 which shows that the integral of the kernel function may be very efficiently approximated by a uniform Riemann sum with inverse step size additive in $\|L\|_{L^{1}}$ and $\log\frac{1}{\epsilon}$. In Section 5, we present an efficient quantum circuit compilation of the LCHS subroutines that proves Theorem 4. In Section 6, we find best-possible LCHS formulas based on the family of kernel functions Equation 6 by numerically optimizing Theorem 1. In Section 7, we prove our new lower bounds and numerically compute best-possible lower bounds based on the QSVT approach.

### v2

@@ 1 Introduction
Non-unitary dynamics appears in other exciting application areas, including dissipative nonlinear dynamics by Carleman linearization [LøKK+21, LAF+23], partial differential equations [CLO21], and linear matrix equations [SLBB25], but exploration is hindered by a more limited understanding of how to best approach Equation˜1. Within the context of quantum algorithms, the goal is to prepare a quantum state that approximates the normalized $\lvert\vec{u}(t)\rangle\doteq\vec{u}(t)/|\vec{u}(t)|$ for some $t>0$, which has the formal solution
@@
in terms of the time-ordering operator $\mathcal{T}$. In the special case of anti-Hermitian $A(t)$ and a zero forcing term $\vec{b}(t)=0$, Equation˜1 is the famous Schrödinger equation that underlies quantum mechanics and the Hamiltonian simulation problem. In the quantum query setting, Equation˜2 is solved by queries to a unitary $\textsc{Prep}_{\lvert\vec{u}_{0}\rangle}$ that prepares the initial state $\lvert\vec{u}_{0}\rangle$ from a computational basis state, and a unitary Definition˜1 that block-encodes the operator $A$ (see Definition˜15 for the time-dependent case), which is a well-established access model that generalizes sparse matrices with efficiently computable entries [ATS03].
@@
[BC24] Dyson series+QLSP $A+A^{\dagger}\succeq 0$ $u_{\text{r}}\alpha_{A}t\log\frac{\alpha_{A}t}{\epsilon}\log\frac{1}{\epsilon}$ $u_{\text{r}}\alpha_{A}t\log\frac{1}{\epsilon}$
@@
In Table˜1, we compare quantum algorithms for the general case of non-Hermitian $A$ with the results of this work. To compare algorithmic advances, we focus on the homogeneous time-independent case where $\vec{u}(t)=e^{-At}\lvert\vec{u}_{0}\rangle$. These algorithms require some ‘stability’ condition on $A$, such as non-negative eigenvalues $\mathrm{Re}[\Lambda]\succeq 0$ [Ber14, BCOW17, CL20, LS24a] in the eigendecomposition $A=S\Lambda S^{-1}$. In this setting, all known results are suboptimal by some logarithmic factor of time $t$, error $\epsilon$, or condition number $\kappa=\|S\|\|S^{-1}\|$, which is a measure of matrix non-normality. Under the stronger condition where A-A has non-positive logarithmic norm or $L\doteq\frac{1}{2}(A+A^{\dagger})\succeq 0$ [Kro23], the recent Linear-Combination-of-Hamiltonian (LCHS) [ALL23, ACL23] approach achieves better scaling, in particular optimal scaling in $u_{\text{lchs}}\doteq|\vec{u}_{0}|/|\vec{u}(t)|$ queries to $\textsc{Prep}_{\lvert\vec{u}_{0}\rangle}$. However, the query complexity to $\textsc{Be}[A/\alpha_{A}]$ scales as $\mathcal{O}(\alpha_{A}t\log^{1+\mathcal{O}(1)}\frac{1}{\epsilon})$, which is suboptimal in error. Furthermore, a no-go result [ACL23, Proposition 7] suggested that achieving the optimal scaling $\mathcal{O}(\alpha_{A}t\log\frac{1}{\epsilon})$ was impossible within the existing LCHS framework. Moreover, known lower bounds are additive in each parameter, in contrast to the multiplicative upper bound. We note that scaling with other matrix- or state-dependent parameters is possible, but these alternate approaches [Kro23, FLT23, LS24a] are incomparable and out-of-scope for this work.
@@
under some additional conditions on $\hat{f}$. Hence, one block-encodes the non-unitary time-evolution operator $\textsc{Be}[e^{-At}/\alpha_{\hat{f}}$] to additive error $\epsilon$ with normalization $\alpha_{\hat{f}}\doteq\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}|\hat{f}(k)|\mathrm{d}k$ by the Linear-Combination-of-Unitaries (LCU) lemma [CW12] after truncating the integral Equation˜4 to finite $R$ and discretizing it to a finite sum. The query complexity is determined by that of Hamiltonian simulation of the ‘most expensive’ unitary $e^{-i(\pm RL+H)t}$. As optimal Hamiltonian simulation in the large time limit requires $\alpha t+\mathcal{O}(\log\frac{1}{\epsilon})$ queries to the block-encoded Hamiltonian [LC17b, LC19, BMPW24], this has query complexity $\mathcal{O}(\alpha_{A}Rt+\log\frac{1}{\epsilon})$. If preparing $\lvert u(t)\rangle$ is desired, this is further multiplied by at least $\alpha_{\hat{f}}$ to account for amplitude amplification. Previously, a LCHS based on $\hat{f}_{\text{ACL23}}(k)\propto\big{(}(1-ik)e^{(1+ik)^{\varphi}}\big{)}^{-1}$ for any constant $\varphi\in(0,1)$ was shown to achieve a truncation threshold $R=\mathcal{O}(\log^{1/\varphi}\frac{1}{\epsilon})$ [ACL23]. Unfortunately, the optimal scaling of $\alpha_{\hat{f}}R=\mathcal{O}(\log\frac{1}{\epsilon})$ was unattainable as $\lim_{\varphi\rightarrow 1}\alpha_{\hat{f}_{\text{ACL23}}}=\infty$, and a no-go theorem [ACL23] forbids this scaling for any LCHS that satisfies Equation˜4.
In this work, we resolve these limitations and present the first query-optimal algorithm for this problem. Our key contributions are: 1) A generalized LCHS framework that achieves the optimal query complexity $\alpha_{\hat{f}}R=\mathcal{O}(\log\frac{1}{\epsilon})$ of block-encoding $e^{-At}$, circumventing a no-go result that applied to previous LCHS methods, and as shown in Figure˜1, with a significant benefit in practical applications; 2) A novel, analytically simple, family of kernel functions that only approximate exponential decay. This is enabled by the generalized LCHS framework and is a conceptual departure from prior work that required an exact match; 3) A greatly simplified and more efficient implementation of the LCHS by proving that a simple, uniform discretization of the LCHS integral converges exponentially fast, leading to simpler quantum circuits than previous non-uniform quadrature methods; 4) New provable and empirical lower bounds on the query complexity, showing that our algorithm’s performance is optimal up to a small constant factor (which we estimate to be approximately three), making further significant improvements in this framework infeasible. These insights are formalized in our main results, which we now summarize.
@@
This is formalized in our main mathematical result Theorem˜1, stated for general time-dependent $A$, which establishes rigorous error bounds on Equation˜5. A consequence of the exact exponential decay equality Equation˜4 is that previous approaches required $\hat{f}(z)$ to decay uniformly to zero in the lower complex plane [ALL23, ACL23], which is also a key assumption of the previous no-go. In contrast, we only require decay $\lim_{|z|\rightarrow\infty}\hat{f}(z)=0$ in an infinite strip of finite width, which implies that $f(x)$ only approximates exponential decay. This flexibility allows us to study a new $4$-parameter family of kernel functions
@@
We prove the optimal cost scaling in our main algorithmic result Theorem˜2 by instantiating Theorem˜1 with the explicit choice $\hat{f}_{2}(k;\gamma,c)\doteq\hat{f}_{2,1}(k;\gamma,c)$ illustrated in Figure˜2 and alternatively defined in terms of elementary functions as follows:
@@
where $\mathrm{erfc}$ is the complementary error function, $c$ is any positive constant, $\gamma=\mathcal{O}(\frac{1}{c}\log^{1/2}\frac{1}{\epsilon})$, and $R=\mathcal{O}(c\gamma^{2})$. We design $f_{2}$ to be analytic in the entire complex plane, which is motivated by the super-polynomial to super-exponential decay of entire functions in the Fourier domain. $f_{2}$ is also the convolution of $e^{-|x+c|+c}$ with a Gaussian kernel, and its simple Fourier transform and constant normalization satisfy the conditions of Theorem˜1 and allow us to derive simple and reasonably tight upper bounds on the performance of our LCHS. Although infinitely many kernel functions satisfy Theorem˜1, not all lead to an optimal-scaling LCHS, even among those where $f$ is entire like $f_{2}$. For instance, the approximations $f_{\text{HMS22}}=\frac{1}{2}e^{-x}\mathrm{erfc}(\gamma-x)$ [HMS+22] and $f_{1}(x;\gamma)$ both turn out to have unbounded norm $\alpha_{\hat{f}}\rightarrow\infty$ when $\gamma\rightarrow\infty$.
Figure 2: Our entire approximation $f_{2}$ to exponential decay converges to $\lim_{\gamma\rightarrow\infty}f_{2}(k,\gamma,c)=e^{-|x+c|+c}$.
@@
For any $t\geq 0$, let $A=L+iH$ be as described in Theorem˜1. For any $\epsilon_{\text{lchs}}\in(0,0.9027$...] and any $c>0$, the LCHS of time-evolution generated by $kL+H$ and the kernel function $\hat{f}_{2}\doteq\hat{f}_{2,1}$ from  Equation˜6 with $\gamma=\frac{1}{c}\sqrt{c+\log\frac{1+1/({2\pi})}{\epsilon_{\text{lchs}}}}$ and $R=2c\gamma^{2}$ satisfies
@@
We also greatly improve the practical implementation of block-encoding the LCHS. We prove in Theorem˜3 that the error $\epsilon_{\text{quad}}$ of a uniform quadrature is exponentially convergent in the stepsize $h$ [TW14]. In contrast, previous analyses either showed polynomial error scaling for the uniform quadrature [ALL23], or used a complicated multi-step non-uniform Gauss-Legendre quadrature [ACL23] with still an asymptotically worse stepsize, possibly a larger block-encoding normalization with no simple closed-form expression due to non-uniform weights, a more complicated quantum circuit implementation for non-uniform quadrature points, and a complicated analysis involving higher derivatives.
@@
Choose the LCHS parameters $\epsilon_{\text{lchs}},R,\gamma,c$ according to Theorem˜2. For any $\epsilon_{\text{quad}}\in(0,4/15]$, and any stepsize ${h}\leq{\pi}/\left(\frac{\|L\|_{L^{1}}}{2}+\log\left(\frac{64e^{3c/2}}{15\epsilon_{\text{quad}}}\right)\right)$ such that $R/h$ is an integer, let $I_{h}$ be the uniform discretization of $O_{R,\gamma,c}(t)$. Then
@@
We then block-encode non-unitary time-evolution in Theorem˜4 to error $\epsilon$ by an explicit and efficient quantum circuit implementation of the Linear-Combination-of-Unitaries (LCU) method [CW12] applied to Theorems˜2 and 3, for instance, with $\epsilon_{\text{quad}}=\epsilon_{\text{lchs}}=\epsilon/3$. Note that for any reasonably small error, the difference in block-encoding normalization $\alpha_{\hat{f}_{2},R,h}$ from $\alpha_{\hat{f}_{2}}$ is extremely small and contributes a multiplicative constant cost that is $1+\mathcal{O}(\epsilon)$. Our use of the uniform quadrature greatly simplifies the explicit quantum circuit implementation, which we provide. As is known in prior art, the query complexity of $\textsc{Sel}_{R}$ for selecting time-evolution for different $k_{j}$ in Theorem˜4, even with the additional control by $\lvert j\rangle$, can be reduced to that of $U_{0}(\pm Rt)$ without any controls by any choice of Hamiltonian simulation algorithm. In the time-independent case, Hamiltonian simulation by quantum signal processing and qubitization [LC17b, LC19] leads to an overall $\mathcal{O}(R\alpha t+\log\frac{1}{\epsilon})=\mathcal{O}(\alpha t\log\frac{1}{\epsilon})$ query complexity.
@@
For any $\epsilon\in(0,4/5]$, and under the same condition as Theorems˜2 and 3, we may block-encode an operator $I_{h}^{\prime}$ such that $\left\|U_{0}(t)-I_{h}^{\prime}\right\|\leq\epsilon$ by
@@
where Prep and $\overline{\textsc{Prep}}$ cost $\mathcal{O}\left(\log(\|L\|_{L^{1}}+\log\frac{1}{\epsilon})\log^{5/2}\frac{1}{\epsilon}\right)$ two-qubit gates, $\textsc{Sel}_{R}$ has the same query complexity $Q$ and gate complexity as simulating $\mathcal{T}e^{-i\int_{0}^{t}\pm RL(s)+H(s)\mathrm{d}s}$ plus $\mathcal{O}\big{(}Q\log(\|L\|_{L^{1}}+\log\frac{1}{\epsilon})\big{)}$ two-qubit gates.
We also optimize the constant factors in the query complexity of our optimal LCHS. The query complexity proportionality constant ${\alpha_{\hat{f}}}\cdot R$ is loosely bounded by Theorem˜2, such as by setting $c=1$ in Figure˜1. Due to our use of the uniform quadrature, the LCU normalization factor $\alpha_{\hat{f},R}\approx\alpha_{\hat{f}}$ up to a vanishingly small error, unlike prior art [PJKW25] where $\alpha_{\hat{f},R}$ can be a significant constant factor larger than $\alpha_{\hat{f}}$ due to non-uniform quadrature weights. We obtain tighter upper bounds on approximation error by evaluating the integrals in Theorem˜1 numerically, and then numerically minimizing the product ${\alpha_{\hat{f}_{2},R}}\cdot R$ for each target $\epsilon$ with respect to $c,\gamma,R$. As seen in Figure˜1, the LCHS with the kernel function $\hat{f}_{2}$ strictly outperforms prior art [ACL23, PJKW25] in all parameter regimes, and further improvements of $\approx 25\%$ are obtained by also optimizing over $j,y$ in the more general family Equation˜6. Moreover, the proofs of implementation details in Theorems˜3 and 4 for the special case $\hat{f}_{2}$ generalize readily to $\hat{f}_{j,y}$.
@@
As $\alpha_{\hat{f}_{2}}=\Theta(1)$ is constant for any constant $c$, the query complexity of our LCHS to state preparation $\textsc{Prep}_{\lvert\vec{u}\rangle_{0}}$, similar to prior LCHS approaches, is optimal due to an $\Omega(u_{\text{lchs}})$ lower bound [ALWZ25, Corollary 8] based on quantum state discrimination. We note that existing lower bounds on $Q_{A}$ are quite loose. When $A$ is anti-Hermitian, this reduces to Hamiltonian simulation with $u_{\text{lchs}}=1$ and $Q_{A}=\Omega(\alpha_{A}t+\frac{\log 1/\epsilon}{\log\log 1/\epsilon})$ [BCK15]. For the general case of $A$ with $L\succeq 0$, it is only known that $Q_{A}=\Omega(\min((\alpha_{A}t)^{\alpha},\log^{\alpha}\frac{1}{\epsilon}))$ for some unknown parameter $\alpha\leq 1$ [ALWZ25]. In Theorem˜5, we present a significantly tighter quantum search lower bound $Q_{A}=\Omega(u_{\text{lchs}}\alpha_{A}t)$, which proves optimality of our LCHS up to a logarithmic factor in $u_{\text{lchs}}$ when $\epsilon$ is constant.
@@
Let $p_{n}$ be the polynomial of degree $n=\min_{f}\widetilde{\mathrm{deg}}_{\epsilon^{\prime}}(f)$, minimized over any $f$ that satisfies Equation˜16 for some constant $\alpha$ – for example, see Figure˜3. By Quantum Singular Value Transformations (QSVT) [GSLW19], $f(A/\alpha_{A})$ is block-encoded with a normalization factor at most $2(1+\epsilon^{\prime})$ using exactly $\widetilde{\mathrm{deg}}_{\epsilon^{\prime}}(f)$ queries to $\textsc{Be}[A/\alpha_{A}]$, which is optimal including constant factors by a matching $\widetilde{\mathrm{deg}}_{\epsilon^{\prime}}(f)$ [MS24, Lan25] lower bound. Hence, our LCHS can be shown to be asymptotically optimal in all parameters if $\min_{f}\widetilde{\mathrm{deg}}_{\epsilon^{\prime}}(f)=\Omega(\tau\log\frac{1}{\epsilon})$, where minimization is over all $f$ satisfying Equation˜16. We compute $p_{n}$ by writing the optimization for $\epsilon$ over $p_{n}$ satisfying Equation˜16 as a semi-infinite linear program [HK93] and in Figure˜3 observe an empirical scaling of $\min_{f}\widetilde{\mathrm{deg}}_{\epsilon^{\prime}}(f)\approx 0.31\tau\log\frac{e}{31\epsilon}$ when $\alpha=e$. To our knowledge, this is the first computation of the optimal bounded polynomial for $e^{-\tau x}$ for $\tau x\geq 0$ and is of independent interest. Based on these empirical fits, our optimized LCHS $\hat{f}_{2}$ and $\hat{f}_{j,y}$ cannot be improved by more than a factor of $4$ and $3$ respectively.
@@
The rest of the manuscript is organized as follows. In Section˜1.1, we discuss some possible extensions of our results and conclude with some open problems. In Section˜2, we prove our main mathematical result Theorem˜1 that a LCHS can be constructed from a kernel function that only approximates exponential decay. In Section˜3, we prove our main algorithmic result Theorem˜2 that a LCHS based on the kernel function $\hat{f}_{2}$ block-encodes $U_{0}(t)$ with the desired scaling in normalization and truncation radius of $\alpha_{\hat{f}_{2}}R=\mathcal{O}(\log\frac{1}{\epsilon})$. In Section˜4, we prove Theorem˜3 which shows that the integral of the kernel function may be very efficiently approximated by a uniform Riemann sum with inverse step size additive in $\|L\|_{L^{1}}$ and $\log\frac{1}{\epsilon}$. In Section˜5, we present an efficient quantum circuit compilation of the LCHS subroutines that proves Theorem˜4. In Section˜6, we find best-possible LCHS formulas based on the family of kernel functions Equation˜6 by numerically optimizing Theorem˜1. In Section˜7, we prove our new lower bounds and numerically compute best-possible lower bounds based on the QSVT approach.